OpenStudy (anonymous):

\[\int\limits_{}^{}\ln(2x+1)dx = xln(2x+1) - \int\limits_{}^{} 2dx/(2x+1) \] u = 2x+1 du = 2dx thus \[xln(2x+1) - \int\limits_{}^{}du/u = xln(2x+1) - \ln|2x+1| + c\] Is this correct?

OpenStudy (turingtest):

not quite, I think we need partial fractions for the second part you dropped an x

OpenStudy (anonymous):

oh yeah you are right I did

OpenStudy (anonymous):

should be \[xln(2x+1) - \int\limits_{}^{} 2xdx/(2x + 1)\]

OpenStudy (turingtest):

better yet, write it as\[\int\ln(2x+1)dx = xln(2x+1) - \int \frac{2xdx}{2x+1}\]however you want to do this I suggest a u-sub\[u=2x+1\implies2x=u-1\]

OpenStudy (anonymous):

u = 2x + 1 du = 2dx thus, x ln(2x+1) - \[\int\limits_{}^{} dux/u\]

OpenStudy (anonymous):

but yeah that isn't right :\

OpenStudy (anonymous):

oh nvm :) \[\ln(2x+1)x - \int\limits_{}^{} ((1-u/2)/u)du\]

OpenStudy (turingtest):

yeah that looks better

OpenStudy (turingtest):

\[u=2x+1\implies2x=u-1\]\[du=2dx\]so we have the second integral as\[\frac12\int1-u^{-1}du\]I think you dropped a 1/2, or at least forgot to put it outside the parentheses

OpenStudy (anonymous):

= \[\ln(2x+1)x - \int\limits_{}^{} ((1-u)/2u)du = \ln(2x+1)x - \int\limits_{}^{}(1/2u)du - \int\limits_{}^{}(u/2u)du\] ok I think I got it

OpenStudy (anonymous):

yeah you are right i did

OpenStudy (turingtest):

simplify, and be careful with your signs; you forgot to distribute a negative\[\ln(2x+1)x - \int\limits_{}^{} ((1-u)/2u)du = \ln(2x+1)x - \int\limits_{}^{}(1/2u)du + \int(\cancel{u}/2\cancel{u}^1)du\]\[\ln(2x+1)x - \int\limits_{}^{} ((1-u)/2u)du = \ln(2x+1)x - \int\limits_{}^{}(1/2u)du -+\int\limits_{}^{}(1/2)du\]

OpenStudy (turingtest):

\[x\ln(2x+1)-\frac12\int1-u^{-1}du\]is the way I prefer to see it :)

OpenStudy (anonymous):

the answer being xln(2x+1) +x/2 -(1/4)ln|2x+1| + c

OpenStudy (anonymous):

crud I made a mistake :l

OpenStudy (turingtest):

\[\frac12\int1-u^{-1}du=\frac12u-\frac12\ln|u|+C=\frac12(2x+1)-\frac12\ln(2x+1)+C\]so we have\[x\ln(2x+1)-\frac12(2x+1)+\frac12\ln(2x+1)+C\]which you can simplify

OpenStudy (turingtest):

\[\ln(2x+1)(x+\frac12)-x-\frac12+C\]since 1/2 is a constant we can absorb it inot C and get\[\ln(2x+1)(x+\frac12)-x+C\]

OpenStudy (anonymous):

I will try it agin Thanks for the help

OpenStudy (turingtest):

welcome :)